OpenStudy (anonymous):

HELP ME PLZZZ!!<3 in this problem determine which equation best represents the problem. The admission fee at a small fair is $1.50 for children and $4.00 for adults. On a certain day, 2200 people enter the fair and $5050 is collected. how many children and how many adults attended? which set of equations would you use to set up this problem? A:4a+c=2200 a+1.5c=5050 B: a+1.5c=5050 4a+c=5050 C: a+c=5050 4a+c=2200 D: a+c=2200 4a+1.5c=5050

OpenStudy (anonymous):

@Ziferot

OpenStudy (texaschic101):

ok....we have adults and children entering....there a total of 2200 people entering a + c = 2200 child tickets cost 1.50 and adult tickets cost 4.....total collected was 5050 1.5c + 4a = 5050 -- can also be 4a + 1.5c = 5050 when doing problems such as these, be sure you do not get the money mixed up with the attendance.

OpenStudy (anonymous):

D

OpenStudy (texaschic101):

yep ^^ thats it.....D

OpenStudy (anonymous):

Thanks guys <3

OpenStudy (texaschic101):

sure thing :)